# Quantum Data Management in the NISQ Era: Extended Version

### v2

@@ 4.2.1. Quantum-specific optimizations
Existing simulators for quantum computing are mostly in Python (cuQuantum team, 2024; Vincent et al., 2022), C++ (AI, 2021; Vincent et al., 2022) or Julia (Luo et al., 2020). There are also dedicated quantum programming languages and domain-specific compilation techniques. See the survey (Heim et al., 2020). In addition, tensor network rewriting techniques have been considered in the ZX-calculus (Coecke and Duncan, 2011). ZX-calculus is a graphical language for representing and reasoning about quantum computations, represented as tensor networks with specialized tensors called spiders (van de Wetering, 2020b). Its graphical nature enables simplifications by locally merging connected vertices while preserving the original tensor, useful in various contexts including formal verification (van de Wetering, 2020b; Lehmann et al., 2022; de Beaudrap and Horsman, 2020).
Quantum-specific optimizations share some conceptual similarities with classical database optimization, yet they diverge significantly due to the unique characteristics of quantum data and operations. First, classical databases rely on relational or NoSQL data models (graph, document-based, etc.) with indexing mechanisms like B-trees or hash tables to facilitate efficient access and retrieval. In simulations, however, quantum states are represented as state vectors (Fig. 2a) or by using more advanced data structures such as tensor networks (Fig. 2b). Second, in classical databases, query optimization involves techniques like join ordering to identify the optimal ordering of join operations between relations for an efficient query plan (Steinbrunn et al., 1997; Leis et al., 2015; Trummer and Koch, 2017). For tensor-based simulators, a defining feature is the optimization of contraction order, which determines the most efficient sequence to contract tensors representing quantum states (Young et al., 2023; cuQuantum team, 2024). Contraction order optimization parallels the role of join order optimization in traditional databases but is more complex due to the exponential growth of quantum state dimensions and the unique features of quantum data, such as entanglement. Third, classical database systems rely heavily on cost estimation to predict and minimize resource use for query execution. Quantum computation simulators, however, lack cost estimation frameworks and instead use optimizations like parallelization, SIMD (Single-Instruction, Multiple-Data) processing, and matrix decomposition techniques (e.g., SVD) to improve performance (Young et al., 2023). Quantum-specific issues, such as quantum noise (Sec. 2), complicate simulation further. These differences lead to the need for fundamentally new optimization frameworks to manage and simulate quantum computation effectively, beyond classical structures and processes. The following question now arises.
@@ 4.2.2. Quantum-specific data representations
Simulating quantum circuits requires precise and efficient representations of quantum states and operations, accounting for the complex nature of quantum mechanics. Ensuring accuracy and scalability as the system grows in complexity is crucial. We mentioned tensor networks as a way to represent the simulation, but various other data representations exist (Vinkhuijzen et al., 2024), e.g., based on algebraic decision diagrams, matrix product states, just to name a few.
@@
An operation on a data structure is tractable if it executes in polynomial time with respect to the input size (Darwiche and Marquis, 2002). The choice of data representation can significantly impact whether certain quantum operations (e.g., gates, measurement explained in Sec. 2 and Appendix A of our technical report (Hai et al., 2025)) are tractable (Vinkhuijzen et al., 2024). We emphasize three main requirements for data representation: expressiveness, closure, and succinctness/tractability.
For expressiveness, the data representation must be sufficiently rich to represent quantum states and measurements. Closure refers to the property that the result of quantum operations on the represented states can also be represented within the same data representation. This property ensures compositionality. Finally, it is crucial to minimize the space required to represent a quantum state, as the number of qubits and the level of entanglement can result in exponentially large states. Therefore, we seek data representations that are succinct, minimizing the space needed to store data while still allowing for efficient query processing. It may also be interesting to consider these questions beyond tensor network representations as well (Bravyi and Gosset, 2016; van de Wetering, 2020a; Zulehner and Wille, 2018; Jozsa and Miyake, 2008). A concrete idea for a novel data representation is to combine MPS and the recently-proposed local-invertible map decision diagram (LIMDD) (Vinkhuijzen et al., 2023). LIMDD compresses a state vector by lumping together parts of the vector that are equivalent modulo simple quantum gates. In polynomial time and space in the number of qubits, LIMDDs can simulate circuits that MPS cannot and vice versa (Vinkhuijzen et al., 2024). One approach to combine these strengths in a single data structure is applying the lumping to the MPS matrices instead of quantum state vectors, by building upon existing work on equivalence characterizations of MPS (Lami and Collura, 2024).
@@ 4.3. More data management opportunities
ML & Quantum data management. i) Simulation: as explained in Sec. 4.2.1, a key challenge in optimizing tensor network based simulation is to find optimal tensor contraction order. A recent research direction is to apply machine learning (ML), specifically reinforcement learning (RL) and graph neural networks (GNN), to optimize tensor contraction order, addressing the computationally intensive nature of this challenge (Meirom et al., 2022; Liu and Zhang, 2023). ii) Error correction: decoding methods for quantum error correction (QEC), like MWPM, face scalability challenges in large quantum systems, where rapid error detection within strict time limits is essential (Vittal et al., 2023). An interesting new direction is data-driven QEC, which employs ML techniques to quantum error correction, such as RL (Zeng et al., 2023; Nautrup et al., 2019; Andreasson et al., 2019), multilayer perceptrons (MLP) (Chu et al., 2022), convolutional neural networks (Chamberland et al., 2023), and GNN (Lange et al., 2023). iii) Improving quantum algorithm efficiency: Quantum algorithms are represented and implemented as quantum circuits, where efficiency can be improved by reducing costly gates like SWAP gate or by minimizing circuit depth and gate count. ML methods, specifically RL (Fan et al., 2022; Fösel et al., 2021) and MLP (Paler et al., 2023), have shown potential in optimizing circuit design to enhance the efficiency of quantum circuits on real devices.
@@ A.4. Other representations
The ZX-calculus (van de Wetering, 2020a) is a graphical language, inspired by tensor networks. In the ZX-calculus, a quantum circuit containing gates from some finite-sized elementary gate set is represented by a graph with a node for each gate; each node is labeled with the type of gate it represents. The ZX-calculus also contains a set of rewrite rules, which for example enable one to prove that two circuits are equivalent. For quantum-circuit simulation specifically, the typical use is to ask for a specific entry in the quantum state vector that the ZX diagram represents. The entry is found by attaching a ZX diagram for the input state to the diagram for the circuit, followed by a sequence of rewrite rules and a decomposition which expresses the entry as a linear combination of smaller diagrams which are easier to directly evaluate.
Finally, there are representations of quantum states other than the ones mentioned in the main text, such as the extended stabilizer (Bravyi and Gosset, 2016) and matchgate (Mocherla et al., 2023) formalisms.
@@ B.3. Experiment setting
Input circuit generation. The inputs of a classical simulator are quantum circuits. In preliminary experiments, we evaluated the baselines and RDBMS solutions over three widely used circuits: W state preparation circuit, GHZ state preparation circuit and QFT circuit. We have developed a circuit generator which is made publicly available online. [^6] https://github.com/infiniData-Lab/Quantum The input circuits are generated in JSON-format with a given number of qubits and a list of gates which are JSON-objects themselves containing an identifier like ”H” for the Hadamard gate and a list of the qubits it acts on. The ideal contraction path is part of the input and generated by the Python library opt_einsum 3.3.0777https://pypi.org/project/opt-einsum/ providing the generated Einstein summation notation in the format ”ij,jk-¿ik”. This notation corresponds to a tensor contraction described by the Einstein summation notation $C_{ik}=\sum_{j}A_{ij}B_{jk}$, where $i,j,k$ are the indices of the corresponding tensors $A$, $B$ and $C$.

### v3

@@ 4.2.1. Quantum-specific optimizations
Existing simulators for quantum computing are mostly in Python (cuQuantum Team, 2024; Vincent et al., 2022), C++ (AI, 2021; Vincent et al., 2022) or Julia (Luo et al., 2020). There are also dedicated quantum programming languages and domain-specific compilation techniques. See the survey (Heim et al., 2020). In addition, tensor network rewriting techniques have been considered in the ZX-calculus (Coecke and Duncan, 2011). ZX-calculus is a graphical language for representing and reasoning about quantum computations, represented as tensor networks with specialized tensors called spiders (van de Wetering, 2020). Its graphical nature enables simplifications by locally merging connected vertices while preserving the original tensor, useful in various contexts including formal verification (van de Wetering, 2020; Lehmann et al., 2022; de Beaudrap and Horsman, 2020).
Quantum-specific optimizations share some conceptual similarities with classical database optimization, yet they diverge significantly due to the unique characteristics of quantum data and operations. First, classical databases rely on relational or NoSQL data models (graph, document-based, etc.) with indexing mechanisms like B-trees or hash tables to facilitate efficient access and retrieval. In simulations, however, quantum states are represented as state vectors (Fig. 2a) or by using more advanced data structures such as tensor networks (Fig. 2b). Second, in classical databases, query optimization involves techniques like join ordering to identify the optimal ordering of join operations between relations for an efficient query plan (Steinbrunn et al., 1997; Leis et al., 2015; Trummer and Koch, 2017). For tensor-based simulators, a defining feature is the optimization of contraction order, which determines the most efficient sequence to contract tensors representing quantum states (Young et al., 2023; cuQuantum Team, 2024). Contraction order optimization parallels the role of join order optimization in traditional databases but is more complex due to the exponential growth of quantum state dimensions and the unique features of quantum data, such as entanglement. Third, classical database systems rely heavily on cost estimation to predict and minimize resource use for query execution. Quantum computation simulators, however, lack cost estimation frameworks and instead use optimizations like parallelization, SIMD (Single-Instruction, Multiple-Data) processing, and matrix decomposition techniques (e.g., SVD) to improve performance (Young et al., 2023). Quantum-specific issues, such as quantum noise (Sec. 2), complicate simulation further. These differences lead to the need for fundamentally new optimization frameworks to manage and simulate quantum computation effectively, beyond classical structures and processes.
Q2 summary. Tensor-based database technologies are promising for boosting simulation performance, with efficiency depending on tensor operations, sparsity, memory layouts, and hardware.
The following question now arises.
@@
Q3 summary. Quantum-specific optimizations require a dedicated simulation optimizer to handle tensor contraction, state sparsity, noise levels, and task-specific strategies, while also eliminating redundant computations through equivalence checking.
@@ 4.2.2. Quantum-specific data representations
Simulating quantum circuits requires precise and efficient representations of quantum states and operations, accounting for the complex nature of quantum mechanics. Ensuring accuracy and scalability as the system grows in complexity is crucial. We mentioned tensor networks as a way to represent the simulation, but various other data representations exist (Vinkhuijzen et al., 2024), e.g., based on algebraic decision diagrams.
@@
An operation on a data structure is tractable if it executes in polynomial time with respect to the input size (Darwiche and Marquis, 2002). The choice of data representation can significantly impact whether certain quantum operations (e.g., gates, measurement explained in Sec. 2 and Appendix A of (Hai et al., 2025)) are tractable (Vinkhuijzen et al., 2024). We emphasize three main requirements for data representation: expressiveness, closure, and succinctness/tractability.
For expressiveness, the data representation must be sufficiently rich to represent quantum states and measurements. Closure refers to the property that the result of quantum operations on the represented states can also be represented within the same data representation. This property ensures compositionality. Finally, it is crucial to minimize the space required to represent a quantum state, as the number of qubits and the level of entanglement can result in exponentially large states. Therefore, we seek data representations that are succinct, minimizing the space needed to store data while still allowing for efficient query processing. It may also be interesting to consider these questions beyond tensor network representations as well (Bravyi and Gosset, 2016; van de Wetering, 2020; Zulehner and Wille, 2018; Jozsa and Miyake, 2008). A concrete idea for a novel data representation is to combine MPS and the recently-proposed local-invertible map decision diagram (LIMDD) (Vinkhuijzen et al., 2023). LIMDD compresses a state vector by lumping together parts of the vector that are equivalent modulo simple quantum gates. In polynomial time and space in the number of qubits, LIMDDs can simulate circuits that MPS cannot and vice versa (Vinkhuijzen et al., 2024). One approach to combine these strengths in a single data structure is applying the lumping to the MPS matrices instead of quantum state vectors, by building upon existing work on equivalence characterizations of MPS (Lami and Collura, 2024).
Q4 summary. Effective data representations for quantum simulations require a balance of expressiveness, closure, and succinctness.
@@ 4.3. More Data Management Opportunities
ML & Quantum data management. i) Simulation: as explained in Sec. 4.2.1, a key challenge in optimizing tensor network based simulation is to find optimal tensor contraction order. A recent research direction is to apply machine learning (ML), specifically reinforcement learning (RL) and graph neural networks (GNN), to optimize tensor contraction order, addressing the computationally intensive nature of this challenge (Meirom et al., 2022; Liu and Zhang, 2023). ii) Error correction: decoding methods for quantum error correction like MWPM, face scalability challenges in large quantum systems, where rapid error detection within strict time limits is essential (Vittal et al., 2023). An interesting new direction is data-driven QEC, which employs ML techniques to quantum error correction, such as RL (Zeng et al., 2023; Nautrup et al., 2019; Andreasson et al., 2019), multilayer perceptrons (MLP) (Chu et al., 2022), convolutional neural networks (Chamberland et al., 2023), and GNN (Lange et al., 2023). iii) Improving quantum algorithm efficiency: Quantum algorithms are represented and implemented as quantum circuits, where efficiency can be improved by reducing costly gates like SWAP gate or by minimizing circuit depth and gate count. ML methods, specifically RL (Fan et al., 2022; Fösel et al., 2021) and MLP (Paler et al., 2023), have shown potential in optimizing circuit design to enhance the efficiency of quantum circuits on real devices.
@@ A.4. Other representations
The ZX-calculus (van de Wetering, 2020) is a graphical language, inspired by tensor networks. In the ZX-calculus, a quantum circuit containing gates from some finite-sized elementary gate set is represented by a graph with a node for each gate; each node is labeled with the type of gate it represents. The ZX-calculus also contains a set of rewrite rules, which for example enable one to prove that two circuits are equivalent. For quantum-circuit simulation specifically, the typical use is to ask for a specific entry in the quantum state vector that the ZX diagram represents. The entry is found by attaching a ZX diagram for the input state to the diagram for the circuit, followed by a sequence of rewrite rules and a decomposition which expresses the entry as a linear combination of smaller diagrams which are easier to directly evaluate.
Finally, there are representations of quantum states other than the ones mentioned in the main text, such as the extended stabilizer (Bravyi and Gosset, 2016) and matchgate (Mocherla et al., 2024) formalisms.
@@ B.3. Experiment setting
Why these baselines? The state-of-the-art approach in (Blacher et al., 2023) treats quantum states as order-n tensors and transforms the simulation circuits into a series of SQL queries executed in relational databases. In Sec. B.1, we further refine this RDBMS-based solution by explicitly defining its relational representation. We select Baseline I because it also adopts an order-n tensor model, enabling a direct comparison of whether the sparse relational encoding in an RDBMS delivers any tangible benefits. We choose Baseline II (MPS) because Matrix Product States are widely acknowledged in the quantum computing community as a more sophisticated and frequently used representation for qubit states. Including MPS allows us to evaluate how well the relational representation competes with this state-of-the-art approach.
As an initial effort, our preliminary findings emphasize representation rather than optimization. As discussed in Sec. 4.2.1, existing research prototypes and commercial simulators often employ parallelization, SIMD processing, and matrix decomposition, which substantially influence efficiency. Addressing Q1, i.e., whether it is advantageous to push simulations into databases or whether a database can exploit memory more effectively than established simulators, remains an open avenue for future research. In particular, leveraging native database features such as indexing, caching, and exploring the emerging tensor-based database technologies mentioned in Q2 will require more system-building efforts to determine how database technologies might be enhanced for large-scale quantum circuit simulations.
Input circuit generation. The inputs of a classical simulator are quantum circuits. In preliminary experiments, we evaluated the baselines and RDBMS solutions over three widely used circuits: W state preparation circuit, GHZ state preparation circuit and QFT circuit. We have developed a circuit generator which is made publicly available online. [^5] https://github.com/infiniData-Lab/Quantum The input circuits are generated in JSON-format with a given number of qubits and a list of gates which are JSON-objects themselves containing an identifier like ”H” for the Hadamard gate and a list of the qubits it acts on. The ideal contraction path is part of the input and generated by the Python library opt_einsum 3.3.0666https://pypi.org/project/opt-einsum/ providing the generated Einstein summation notation in the format ”ij,jk-¿ik”. This notation corresponds to a tensor contraction described by the Einstein summation notation $C_{ik}=\sum_{j}A_{ij}B_{jk}$, where $i,j,k$ are the indices of the corresponding tensors $A$, $B$ and $C$.